OpenStudy (cenaida):

Find the slope of the line that contains the following pair of points. (1, 2), (5, 2) A. -1 B. 0 C. 4 D. Undefined is it C?

OpenStudy (cenaida):

@hockeychick23

OpenStudy (hockeychick23):

(1, 2), (5, 2) 2-2/5-1= 0/4= 0

OpenStudy (cenaida):

oh.. ok i get it thanks again!

OpenStudy (anonymous):

its 0!!!!!!

OpenStudy (hockeychick23):

sure!

OpenStudy (zale101):

@cenaida apply the same formula i mentioned on your previous question.

OpenStudy (anonymous):

\[y _{2}-y_{1}/ x _{2}-x _{1} \]

OpenStudy (zale101):

Next time, i suggest working on the problem on your post first so we can see where you're having difficulties with.